# Deep Reinforcement Learning for Traveling Purchaser Problems

### v5

@@ 5.1.3 Decoder
The bipartite graph of the TPP instance $U$ is encoded through the input embedding module and market encoder. After that, the decoder is executed iteratively to construct the route, one node at a step. At each step $t=1,2,...,T$, the decoder takes as input a decoding context $h_{d}$ of the state $s_{t}$, and outputs a distribution $p_{\theta}(a_{t}\mid s_{t})$ over the actions $a_{t}\in A_{t}$. Then, an action $a_{t}$ is selected based on the distribution. The corresponding market/depot is added to the end of the partial route, and the state is updated accordingly to start the next decoding step, until a complete route is constructed.
@@
The demand context $g_{\mathcal{K}}^{t}$ contains the information of remaining demand $d^{t}$. Third, the route context $o_{\pi}^{t}$ is the embedding for the partial route $\pi^{t}$, i.e., a sequence of market/depot nodes. We adopt an LSTM (Hochreiter and Schmidhuber, 1997) recurrent network, similar to the decoding RNN in Ptr-Net (Vinyals et al., 2015), to obtain the route context $o_{\pi}^{t}$:
@@
where $h_{v^{t}}^{N}$ is the final embedding of the node $v^{t}$ added to the partial route at the last step.
The decoder leverages the information from the decoding context $h_{d}$ to make decisions. It first processes the decoding context $h_{d}$ through a one-to-many attention layer, where the query is from the decoding context $h_{d}$, and the keys and values are from the final market/depot node embeddings $h_{0}^{N},\ldots,h_{\left|M\right|}^{N}$. Then, the action distribution $p_{\theta}\left(a_{t}\mid s_{t}\right)$ is computed as the single-head attention weights between the updated decoding context vector (denoted as $h_{d}^{\prime}$) and the final node embeddings $h_{0}^{N},\ldots,h_{\left|M\right|}^{N}$ of markets/depot. We first computed attention scores between $h_{d}^{\prime}$ and the final market/depot node embeddings:
@@
Then, an action $a_{t}$ is selected based on the distribution, and the corresponding market/depot $m$ is added to the partial route. Specifically, we sample the action $a_{t}$ from the distribution $p_{\theta}\left(a_{t}\mid s_{t}\right)$ during training, and select greedily the action with the maximum probability when applying the learned policy for inference.
We remark that besides its strong capability, the design of our policy network has many benefits. First, our policy network based on GNNs and MHA is agnostic to the size of instances being solved, as long as the input feature dimensions are fixed, which is achieved in our proposed bipartite graph representation. This allows the policy network to directly adapt to TPP instances with varying numbers of markets and products, without the need for parameter adjustment or retraining. Second, the embedding process for the static part $U$ and the dynamic part $\pi^{t}$ and $d^{t}$ of a state can be separated, which improves the computational efficiency of the policy network. Specifically, the embeddings of $U$ are pre-computed only once at the initial states using the computationally intensive input embedding module and market encoder, while the computationally light decoder is executed repeatedly with varying decoding contexts that contain the dynamic information of states. Moreover, the forward-propagation of our policy network based on GNNs and MHA is highly parallelizable, which further improves the computational efficiency when used for solving TPPs.
5.2 Basic Training Algorithm
@@
The policy network is updated based on the reward received after the solving process. As introduced in Section 3.2.4 and Section 3.2.5, given a TPP instance $U$, the policy network $\theta$ generates the final route $\pi$ with probability $p_{\theta}(\pi\mid U)=\prod_{t=1}^{T}p_{\theta}(a_{t}\mid s_{t})$ during training, where the associated traveling cost can be directly accessed, and the purchasing cost is obtained by solving a transportation problem. To align with the notations in DRL, we let $L(\pi\mid U)$ denote the loss of $\pi$, which is defined as the negative reward for the terminal state. We update the policy network $\theta$ to minimize the expectation of loss $L(\pi\mid U)$, using the REINFORCE algorithm (Williams, 1992) with baseline $b(U)$:
$\nabla_{\theta}\mathcal{L}(\theta\mid U)=\mathbb{E}_{\pi\sim p_{\theta}(\pi\mid U)}\left[(L(\pi\mid U)-b(U))\nabla_{\theta}\log p_{\theta}(\pi\mid U)\right],$
where the baseline $b(U)$ can be seen as an estimation for the difficulty of instance $U$ to measure the relative advantage of route $\pi$ generated by the policy network. A well-defined baseline can significantly reduce the variance of gradients and accelerate training. We define the baseline $b(U)$ as the loss of the route obtained from a deterministic greedy rollout, i.e., where the action with the highest probability is always selected at each decision step, based on a baseline network $\theta^{\text{BL}}$ (Kool et al., 2019). Similar to the target Q-network in DQN (Mnih et al., 2015), the baseline network $\theta^{\text{BL}}$ is a copy of policy network $\theta$, but is fixed during each epoch to stabilize the baseline value. The baseline network $\theta^{\text{BL}}$ is periodically replaced by the latest policy network $\theta$ if the improvement of the policy is significant enough according to a paired t-test. Based on the gradient $\nabla_{\theta}\mathcal{L}(\theta\mid U)$, the parameters of policy network are updated using the Adam optimizer (Kingma and Ba, 2015). The basic training algorithm introduced above is outlined as Algorithm 1.
@@ 6 Meta-learning Strategy
Despite our proposed framework and the key DRL components introduced above, how to effectively learn a high-quality policy remains another challenge, especially on large-sized TPP instances. First, due to the huge state-action space for solving TPPs, the training suffers heavily from low sample efficiency. The policy network often makes random and suboptimal actions in the early training stage, requiring substantial trial-and-error before a good policy can be found. More seriously, when trained on large-sized TPP instances, a randomly initialized policy network may completely fail to find an effective or even reasonable solution and result in training collapse (see Figure 4(b)). Second, while a well-trained policy can produce high-quality solutions for TPP instances from the same distribution as training, it tends to exhibit poor generalization performance on instances from different distributions. Generalization across different instance distributions is a highly desirable property for learning-based methods for solving CO problems.
@@
Through meta-training, the meta policy network can serve as an effective initialization that achieves good performance on a new instance distribution through a few fine-tuning steps, even on large-sized instances that tend to experience training collapse if the policy network is trained from scratch. Moreover, the meta policy network can learn cross-distribution knowledge through meta-learning, thus generalizing well across varied instance sizes and distributions, even showing good zero-shot generalization performance on instances from a different distribution that is never seen during training.
@@ 7 Experiments
We conduct extensive experiments on various synthetic TPP instances and the TPPLIB benchmark to evaluate our DRL-based approach, comparing it with several well-established TPP heuristics. Additionally, we further confirm the zero-shot generalization ability of our policy network on larger-sized TPP instances that are unseen during training.
@@ 7.1.1 Instances
We first randomly generate synthetic TPP instances of sizes $\left(|M|,|K|\right)=(50,50)$, $(50,100)$, $(100,50)$, and $(100,100)$ for training and evaluation, where $|M|$ is the number of markets and $|K|$ is the number of products. The generation of U-TPP and R-TPP instances follows the standard procedure introduced in (Laporte et al., 2003), corresponding to Class 3 and Class 4 in the TPPLIB benchmark (Riera-Ledesma, 2012). For the U-TPP instance, $(|M|+1)$ integer coordinates (including the depot) are randomly generated within the $[0,1000]\times[0,1000]$ square according to a uniform distribution, and the traveling costs are defined as the truncated Euclidean distances. Each product $k\in K$ is supplied at $|M_{k}|$ randomly selected markets, where $|M_{k}|$ is uniformly generated in $\left[1,|M|\right]$. The price $p_{ik}$ for each product $k\in K$ and each market $i\in M_{k}$ is randomly sampled in $[1,10]$. The R-TPP instances are generated in the similar manner as U-TPP, with an additional limit $q_{ik}$ on the available supply quantities, which is randomly sampled in $[1,15]$ for each product $k\in K$ and each market $i\in M_{k}$. A parameter $\lambda$ is used to control the number of markets in a feasible solution through the product demand $d_{k}:=\left\lceil\lambda\max_{i\in M_{k}}q_{ik}+(1-\lambda)\sum_{i\in M_{k}}q_{ik}\right\rceil$ for $k\in K$, where $10<\lambda<$ 1. We employ $\lambda=\left\{0.99,0.95,0.9\right\}$ in our experiments, corresponding to the most difficult instances in the TPPLIB (Manerba et al., 2017).
We define the U-TPP instances sharing the same $\left(|M|,|K|\right)$, or the R-TPP instances sharing the same $(|M|,|K|,\lambda)$ as an instance distribution. For each instance distribution, training instances are generated on-the-fly, and a set of 1000 instances are generated for evaluation. Moreover, we further generate larger-sized instances for evaluating the zero-shot generalization ability of the policy network, which is trained on smaller instances. In addition to the synthetic instances, the well-known TPPLIB benchmark is also used for evaluation, which includes 5 instances for each instance distribution, called a category.
@@ 7.1.2 Baseline Methods
We select several well-established TPP heuristics reported in (Manerba et al., 2017), including GSH, CAH, and TRH. A widely used and effective practice is to incorporate constructive heuristics (e.g., GSH and CAH) with TRH to remove redundant markets as soon as a solution is produced. Additionally, the solution can potentially be further improved by using a TSP solver to re-sequence the order of visited markets. In our experiment, we employ GSH and CAH, both followed by TRH and the TSP re-sequence, as the baselines, denoted as “GSH+TRH” and “CAH+TRH”, respectively.
Note that we do not include exact methods in the evaluation on synthetic instances since it can take several days to weeks to solve an evaluation instance set using exact methods (even for U-TPP). For a comparison of the optimality gap, we report the best-known solutions in the literature (Riera-Ledesma, 2012) on the TPPLIB benchmark instances to assess the optimality gaps of the baseline heuristics and our DRL-based approach.
@@ 7.1.3 Training and Inference of the Policy Network
In the evaluation of our DRL-based approach, we execute the inference of the learned policy network with greedy decoding (similar to the greedy rollout baseline) and instance augmentation (Kwon et al., 2020). The produced route and purchasing plan can be directly used as an end-to-end solution to the TPP instance, denoted as “RL-E2E”. Similar to the baseline heuristics, this end-to-end solution can be further improved through a post-optimization procedure (TRH and TSP re-sequence), denoted as “RL+TRH”. For a fair comparison, the baseline heuristics and our approach are implemented on the same machine and environment.
@@ 7.2 Training Performance
However, the policy network suffers serious training collapse when trained from scratch on large-sized U-TPP and R-TPP instances of sizes $\left(|M|,|K|\right)=(100,50)$ and $(100,100)$. As shown in Figure 4(b), the policy network trained from scratch on U-TPP instances of size $\left(|M|,|K|\right)=(100,50)$ is stuck in a high-loss region, and the loss completely fails to decrease throughout training. This is mainly because the state-action space is so large that the policy network cannot find an effective or even reasonable solution if the training starts from a random initialization. Therefore, we instead apply our meta-learning strategy to train the policy network for U-TPP and R-TPP instances with $\left(|M|,|K|\right)=(100,50)$ and $(100,100)$. As shown in Figure 4(b), the utilization of meta-learning strategy significantly promotes the training for an effective policy.
Figure 4: The training curve on (a) U-TPP instances of size $\left(|M|,|K|\right)=(50,50)$, and (b) U-TPP instances of size $\left(|M|,|K|\right)=(100,50)$
@@ 7.3 Results on Synthetic Instances
Following the training procedures described above, we train the policy network from scratch for U-TPP and R-TPP instances of sizes $\left(|M|,|K|\right)=(50,50)$ and $(50,100)$, and apply the meta-learning strategy for training U-TPP and R-TPP instances of sizes $\left(|M|,|K|\right)=(100,50)$ and $(100,100)$. Then, the learned policy network is evaluated on the set of 1000 synthetic evaluation instances for each instance distribution. We report the metrics of 1) the average objective value and 2) the runtime (in seconds) for each evaluation instance set. The results are reported in Table 2 for U-TPP and Table 3 for R-TPP.
In terms of the objective value, the performance of baseline heuristics varies across different instance distributions. In comparison, our DRL-based approach consistently outperforms both GSH+TRH and CAH+TRH on all synthetic U-TPP and R-TPP instance sets of different sizes and distributions. Notably, the end-to-end solutions (RL-E2E) can already achieve better objective value than the hand-crafted baseline heuristics, and the post-optimization procedure (RL+TRH) further improves the solution, with only a slight increase in solution time. The results demonstrate that our DRL-based approach is able to learn effective policies for solving TPPs of different sizes and distributions.
In terms of the runtime, all baseline heuristics and our DRL-based approach can produce a solution within an average of 0.1 seconds for U-TPP instances of all four sizes. In general, GSH+TRH takes the least time, but the advantage is approximately negligible in practice. However, for more challenging R-TPP instances, the solution time of baseline heuristics increases significantly with the instance size, as shown in Figure 5(b)-(d). GSH+TRH and CAH+TRH take an average of 1.666 seconds and 2.637 seconds, respectively, to produce a solution for R-TPP instances with $|M|=100,|K|=100,\lambda=0.9$. This is mainly because they need to frequently compute the objective values and savings each time a new market is inserted into the route. In contrast, the actions in our DRL-based approach are entirely based on the forward-propagation of the policy network. Therefore, the solution time is approximately linear with the number of visited markets in the constructed route. On synthetic R-TPP instances, the average solution time of RL-E2E is consistently less than 0.1 seconds, and the average solution time of RL+TRH is consistently less than 0.2 seconds.
@@
Table 2: Results on synthetic U-TPP instances.

### v6

@@ 5.1.3 Decoder
The bipartite graph of the TPP instance $U$ is encoded through the input embedding module and market encoder. After that, the decoder is executed iteratively to construct the route, one node at a step. At each step $t=1,2,...,T$, the decoder takes as input a decoding context $h_{d}$ of the state $s_{t}$, and outputs a distribution $p_{\theta}(a_{t}\mid s_{t})$ over the actions $a_{t}\in A_{t}$. Then, an action $a_{t}$ is selected based on the distribution. The corresponding market or depot is added to the end of the partial route, and the state is updated for the next decoding step, until a complete route is constructed.
@@
The demand context $g_{\mathcal{K}}^{t}$ contains the information of remaining demand $d^{t}$, which helps the decoder consider which products are more inadequate when deciding the following markets.
Third, the route context $o_{\pi}^{t}$ is the embedding for the partial route $\pi^{t}$, i.e., a sequence of market and depot nodes. We adopt an LSTM (Hochreiter and Schmidhuber, 1997) recurrent network, which updates the route context $o_{\pi}^{t}$ as follows:
@@
where $h_{v^{t}}^{N}$ is the final embedding of the node $v^{t}$ added to the partial route at the last step. The LSTM helps capture the information about previously selected nodes and guiding the next decision in the route construction process.
The decoder leverages the information from the decoding context $h_{d}$ to make decisions. It first processes the decoding context $h_{d}$ through a one-to-many attention layer, where the query is from the decoding context $h_{d}$, and the keys and values are from the final market / depot node embeddings $h_{0}^{N},\ldots,h_{\left|M\right|}^{N}$. The action distribution $p_{\theta}\left(a_{t}\mid s_{t}\right)$ is computed as the single-head attention weights between the updated decoding context vector (denoted as $h_{d}^{\prime}$) and the final node embeddings $h_{0}^{N},\ldots,h_{\left|M\right|}^{N}$ of markets / depot. We first computed attention scores between $h_{d}^{\prime}$ and the final market / depot node embeddings:
@@
Then, an action $a_{t}$ is selected based on the distribution, and the corresponding market / depot $m$ is added to the partial route. Specifically, during training, the action is sampled from $p_{\theta}\left(a_{t}\mid s_{t}\right)$, while during inference, the action with the highest probability is selected greedily.
@@
We highlight several advantages of our policy network design, which contribute to its capability, scalability, and efficiency. First, our policy network, based on GNNs and MHA, is agnostic to the size of instances being solved, given that the input feature dimensions are fixed by use of our bipartite graph representation. This enables the policy network to adapt directly to TPP instances with varying numbers of markets and products, without the need of parameter adjustment or retraining. Second, the embedding process for the static part $U$ and the dynamic part $\pi^{t}$ and $d^{t}$ of a state can be separated, which improves the computational efficiency of the policy network. Specifically, the embeddings of $U$ are pre-computed only once using the computationally intensive input embedding module and market encoder, while a computationally light decoder is executed repeatedly with varying decoding contexts that contain the dynamic information of states. This separation allows for efficient reuse of static embeddings and avoids redundant computations. Moreover, the forward-propagation of our policy network based on GNNs and MHA is highly parallelizable, which further improves the computational efficiency when used for solving TPPs.
5.2 Basic Training Algorithm
The policy network is updated based on the reward received after the solving process. As described earlier, given a TPP instance $U$, the policy network $\theta$ generates the final route $\pi$ with probability $p_{\theta}(\pi\mid U)=\prod_{t=1}^{T}p_{\theta}(a_{t}\mid s_{t})$ during training, where the associated traveling cost can be directly accessed, and the purchasing cost is obtained by solving a transportation problem. To align with the notations in DRL, we let $L(\pi\mid U)$ denote the loss of $\pi$, which is defined as the negative reward for the terminal state. We update the policy network $\theta$ to minimize the expectation of loss $L(\pi\mid U)$, using the REINFORCE algorithm (Williams, 1992) with baseline $b(U)$:
$\nabla_{\theta}\mathcal{L}(\theta|U)=\mathbb{E}_{\pi\sim p_{\theta}(\pi|U)}\left[(L(\pi|U)-b(U))\nabla_{\theta}\log p_{\theta}(\pi|U)\right],$
where the baseline $b(U)$ can be seen as an estimation for the difficulty of instance $U$, which measures the relative advantage of route $\pi$ generated by the policy network. A well-defined baseline can significantly reduce the variance of gradients and accelerate training. We define the baseline $b(U)$ as the loss of the route obtained from a deterministic greedy rollout, where the action with the highest probability is always selected at each decision step, based on a baseline network $\theta^{\text{BL}}$ (Kool et al., 2019). Similar to the target Q-network in DQN (Mnih et al., 2015), the baseline network $\theta^{\text{BL}}$ is a copy of policy network $\theta$, but fixed during each epoch to stabilize the baseline value. The baseline network $\theta^{\text{BL}}$ is periodically replaced by the latest policy network $\theta$ if the improvement of the policy is significant enough according to a paired t-test. Based on the gradient $\nabla_{\theta}\mathcal{L}(\theta\mid U)$, the parameters of policy network are updated using the Adam optimizer (Kingma and Ba, 2015). The basic training algorithm introduced above is outlined as Algorithm 1.
@@ 6 Meta-learning Strategy
Despite our proposed framework and the key DRL components, how to effectively learn a high-quality policy remains another challenge, especially for large-sized TPP instances. First, due to the huge state-action space in solving TPPs, the training suffers from low sample efficiency. The policy network often makes random and suboptimal actions in the early training stage, requiring substantial trial-and-error before a good policy can be obtained. This issue is exacerbated when training on large-sized TPP instances, where a randomly initialized policy network may fail to find a reasonable solution, resulting in potential training collapse (see Figure 4(b)). Second, while a well-trained policy can produce high-quality solutions for TPP instances from the same distribution as training, it tends to exhibit poor generalization performance on instances from different distributions. Generalization across different instance distributions is a highly desirable property for learning-based methods for solving CO problems.
@@
Through meta-training, the meta policy network can serve as an effective initialization that achieves good performance on a new instance distribution through a few fine-tuning steps. This is especially beneficial for large-sized instances, which tend to experience training collapse if the policy network is trained from scratch. Moreover, the meta policy network can learn cross-distribution knowledge through meta-learning, enhancing its ability to generalize across varied instance sizes and distributions, even showing good zero-shot generalization performance on instances from a different distribution that is never seen during training.
@@ 7 Experiments
We conduct extensive experiments on various synthetic TPP instances and the TPPLIB benchmark to evaluate our DRL-based approach, comparing it against several well-established TPP heuristics. Furthermore, we validate the zero-shot generalization ability of our policy network on larger-sized TPP instances that are unseen during training. In addition, an ablation study to analyze the contributions of each component in our framework and extended results are presented in the appendices.
@@ 7.1.1 Instances
We first randomly generate synthetic TPP instances of sizes $\left(|M|,|K|\right)=(50,50)$, $(50,100)$, $(100,50)$, and $(100,100)$ for training and evaluation, where $|M|$ is the number of markets and $|K|$ is the number of products. The generation of U-TPP and R-TPP instances follows the standard procedure introduced in (Laporte et al., 2003), corresponding to Class 3 and Class 4 in the TPPLIB benchmark (Riera-Ledesma, 2012). For the U-TPP instance, $(|M|+1)$ integer coordinates (including the depot) are randomly generated within the $[0,1000]\times[0,1000]$ square according to a uniform distribution, and the traveling costs are defined as the truncated Euclidean distances. Each product $k\in K$ is supplied at $|M_{k}|$ randomly selected markets, where $|M_{k}|$ is uniformly generated in $\left[1,|M|\right]$. The price $p_{ik}$ of each product $k\in K$ at each market $i\in M_{k}$ is randomly sampled in $[1,10]$. The R-TPP instances are generated in a similar manner as U-TPP, with an additional limit $q_{ik}$ on the available supply quantities, which is randomly sampled in $[1,15]$ for each product $k\in K$ and each market $i\in M_{k}$. A parameter $\lambda$ is used to control the number of markets in a feasible solution through the product demand $d_{k}:=$ $\left\lceil\lambda\max_{i\in M_{k}}q_{ik}+(1-\lambda)\sum_{i\in M_{k}}q_{ik}\right\rceil,0<\lambda<1$, for $k\in K$. We choose $\lambda=\left\{0.99,0.95,0.9\right\}$ in our experiments, corresponding to the most difficult instances in TPPLIB (Manerba et al., 2017).
We define an instance distribution as U-TPP instances sharing the same $\left(|M|,|K|\right)$, or R-TPP instances sharing the same $(|M|,|K|,\lambda)$. For each instance distribution, training instances are generated on-the-fly, and a collection of 1000 instances is generated for evaluation. Furthermore, we generate larger-sized instances for evaluating the zero-shot generalization ability of the policy network trained on smaller instances. In addition, the well-known TPPLIB benchmark is also used for evaluation, which includes 5 instances for each instance distribution, called a category.
@@ 7.1.2 Baseline Methods
We select several well-established TPP heuristics reported in (Manerba et al., 2017), including GSH, CAH, and TRH. A widely used and effective practice is to incorporate constructive heuristics (e.g., GSH and CAH) with TRH to remove redundant markets as soon as a solution is produced. Additionally, the solution can potentially be further improved by using a TSP solver to re-sequence the visited markets. In our experiments, we implement GSH and CAH, both followed by TRH and the TSP re-sequence, as the baselines, denoted as “GSH+TRH” and “CAH+TRH”, respectively.
Note that we do not include exact methods in the evaluation on synthetic instances, as it can take several days to weeks to solve an evaluation instance set using exact methods (even for U-TPP). For a comparison of the optimality gap, we reference the best-known solutions in the literature (Riera-Ledesma, 2012) on the TPPLIB benchmark instances to compare the optimality gaps of the baseline heuristics and our DRL-based approach.
Table 1: Training configuration
@@ 7.1.3 Training and Inference of the Policy Network
For evaluating our DRL-based approach, we perform inference of the learned policy network using greedy decoding (similar to the greedy rollout baseline) with instance augmentation (Kwon et al., 2020). The produced route and purchasing plan can directly serve as an end-to-end solution to the TPP instance, which we denote as “RL-E2E”. Similar to the baseline heuristics, we can also apply a post-optimization procedure (TRH and TSP re-sequence) to further improve the end-to-end solution, denoted as “RL+TRH”. For a fair comparison, the baseline heuristics and our approach are implemented on the same machine and environment.
@@ 7.2 Training Performance
However, the policy network suffers serious training collapse when trained from scratch on large-sized U-TPP and R-TPP instances of sizes $\left(|M|,|K|\right)=(100,50)$ and $(100,100)$. As shown in Figure 4(b), the policy network trained from scratch on U-TPP instances of size $\left(|M|,|K|\right)=(100,50)$ is stuck in a high-loss region, and the loss completely fails to decrease throughout training. This is mainly because the state-action space is so large that the policy network cannot find a reasonable solution if the training starts from a random initialization. Therefore, we instead apply our meta-learning strategy to train the policy network for U-TPP and R-TPP instances with $\left(|M|,|K|\right)=(100,50)$ and $(100,100)$. As shown in Figure 4(b), the utilization of meta-learning strategy significantly promotes the training for an effective policy.
Figure 4: The training curve on (a) U-TPP instances of size $\left(|M|,|K|\right)=(50,50)$, and (b) U-TPP instances of size $\left(|M|,|K|\right)=(100,50)$.
@@ 7.3 Results on Synthetic Instances
Following the training procedures described above, we train the policy network from scratch for U-TPP and R-TPP instances of sizes $\left(|M|,|K|\right)=(50,50)$ and $(50,100)$, and apply the meta-learning strategy for training U-TPP and R-TPP instances of sizes $\left(|M|,|K|\right)=(100,50)$ and $(100,100)$. Then, the learned policy network is evaluated on the synthetic evaluation instance set for each instance distribution. We report the metrics of 1) the average objective value of the obtained solution and 2) the average runtime to find this solution (in seconds) per instance. The results are reported in Table 2 for U-TPP and Table 3 for R-TPP.
In terms of the objective value, the performance of baseline heuristics varies across different instance distributions. In contrast, our DRL-based approach consistently outperforms both GSH+TRH and CAH+TRH on all synthetic U-TPP and R-TPP instance sets of different sizes and distributions. Notably, the end-to-end solutions (RL-E2E) can already achieve better objective value than the hand-crafted baseline heuristics, and the post-optimization procedure (RL+TRH) further improves the solution, with only a slight increase in runtime. The results demonstrate that our DRL-based approach can learn effective policies for solving TPPs of different sizes and distributions.
Table 2: Results on synthetic U-TPP instances
@@
Table 3: Results on synthetic R-TPP instances
@@
In terms of the runtime, all methods, including the baseline heuristics and our DRL-based approach, can produce a solution within an average of 0.1 seconds for U-TPP instances of all four sizes. Generally, GSH+TRH is slightly faster, but the difference is approximately negligible in practice. However, for more challenging R-TPP instances, the runtime differences become more pronounced. As shown in Figure 5(b)-(d), the runtime of baseline heuristics increases significantly with the instance size. GSH+TRH and CAH+TRH take an average of 1.666 seconds and 2.637 seconds, respectively, to produce a solution for R-TPP instances with $|M|=100,|K|=100,\lambda=0.9$. This is mainly because they need to frequently compute the objective values and savings each time a new market is inserted into the route. In contrast, the actions in our DRL-based approach are entirely based on the forward-propagation of the policy network. Therefore, the solution time is approximately linear with the number of visited markets in the constructed route. On synthetic R-TPP instances, the average solution time of RL-E2E is consistently below 0.1 seconds, and the average solution time of RL+TRH is consistently below 0.2 seconds.